OpenStudy (anonymous):

The length of a rectangle is 6 cm less than twice its width. Find the dimensions of the rectangle if its area is 108 cm^2.

OpenStudy (anonymous):

L=2W-6 108= W* (2W-6) 108 =2W^2 -6W 0 = (2W^2 -6W -108)2 0 = W^2-3W-54 (W+6)(W-9)= 0 W=-6 W=9 is solution L=(2*9)-6 L= 12